# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 3.3 Large chemical potential limit
$\delta\bar{\mu}_{\chi}(t,z)$ $=\frac{|\alpha_{V}|}{\sqrt{3\pi}}\frac{\delta\mu_{V}}{2}\left(e^{-(z-\tilde{v}t)/2\sigma^{2}}-e^{-(z+\tilde{v}t)/2\sigma^{2}}\right),$ (84)
while $\delta\bar{\mu}_{\tilde{\chi}}(t,z)\simeq\delta\bar{\mu}_{\chi}(t,z)$. As shown in Fig. 3, the above solution provides a good match to the numerical solution in the considered limit at the normalized vector chemical potential $\alpha_{V}=\mu_{V}/T=20$. The amplitude of $\delta\bar{\mu}_{\tilde{\chi}}=\delta\bar{\mu}_{A}-s_{V}\delta\bar{\mu}_{H}$ is finite, contrary to the degenerate limit of the case when the charge is conserved when the amplitude of this mode increases like $\alpha_{V}^{4}$ (see Fig. 4 and Eq. (5.19) from Ref. HVWp1 ).
@@
$\tau_{H}^{-1}(\mu_{V}\gg T)=\frac{\tau^{-1}_{\textrm{deg.}}}{\alpha_{V}}\left(1+O\left(\frac{1}{\alpha_{V}}\right)\right)e^{-\alpha_{V}}+O\left(e^{-2\alpha_{V}}\right)\,,$ (85)
with $\tau_{\textrm{deg.}}$ constant. Taking this dependence into account, we can take the degenerate limit and the helicity conservation limit in a correlated manner. While the propagating modes $\omega_{\pm}$ remain unaffected, the non-propagating mode $\omega_{\textrm{im.}}$ becomes, to leading order,
$\omega_{\textrm{im.}}=-\dfrac{i\pi^{2}}{24\tau_{\textrm{deg.}}\alpha_{V}}\,.$ (86)
Therefore, the dissipation is inhibited as we increase the chemical potential, contrary to the naive $\alpha_{V}\to\infty$ limit in Eq. (79). Finally, taking the limit for the amplitudes (82) gives
$\delta\mu_{\chi}^{\textrm{im.}}\simeq-\dfrac{2}{\alpha_{V}^{2}}e^{-|\alpha_{V}|}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}\,\qquad\delta\mu_{V}^{\textrm{im.}}\simeq-\dfrac{72i\kappa_{\Omega}\tau_{\textrm{deg}}}{\pi^{2}\alpha_{V}}e^{-|\alpha_{V}|}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}$ (87)
@@ 3.4 Generic properties of spectra
In Subsec. 3.1 we have seen that, at small vector chemical potential, there is a critical wavenumber above which two of the three modes start to propagate. At the same time, for small wavenumbers at finite chemical potential, we showed in Subsec. 3.2 that again, there are two modes that always propagate. Both results are compatible with each other only if, for small chemical potential, the modes that propagate at small $k$ turn into non-propagating modes at some finite $k$. We will shortly see that this is indeed the case.
Figure 4: Propagating points $\kappa_{\Omega}^{+}$ (upper black curve) and $\kappa_{\Omega}^{-}$ (lower black curve) as functions of the normalized vector chemical potential $\alpha_{V}=\mu_{V}/T$ for the unpolarized plasma. In the shaded region, there are no propagating modes. The boundaries $\kappa_{\Omega}=\kappa_{\Omega}^{\pm}$ of the shaded region, defined in Eq. (93), join each other at the critical value $\kappa_{\Omega}^{\rm crit.}$ given in Eq. (98). The vertical lines mark the energy of the collision $\sqrt{s}$ corresponding to the given value of the normalized chemical potential $\alpha_{V}$ [see a discussion around Eq. (99).]
@@
$04ac^{3}(\kappa_{\Omega}\tau_{H})^{6}-(b^{2}c^{2}+18abcd-27a^{2}d^{2})(\kappa_{\Omega}\tau_{H})^{4}+4b^{3}d(\tau_{H}\kappa_{\Omega})^{2}=$ 0 . (92)
@@
$(\kappa_{\Omega}^{\pm})^{2}=\frac{1}{8ac^{3}\tau_{H}^{2}}\left[b^{2}c^{2}+18abcd-27a^{2}d^{2}\pm\sqrt{bc-ad}(bc-9ad)^{3/2}\right].$ (93)
@@
$\kappa_{\Omega}^{-}=\dfrac{\pi^{2}}{21\tau_{H}\ln^{2}2}\alpha_{V}+O(\alpha_{V}^{3}),\quad\kappa_{\Omega}^{+}=\frac{\pi^{2}}{12\tau_{H}\ln 2}+\dfrac{1.135}{\tau_{H}}\alpha_{V}^{2}+O(\alpha_{V}^{4}).$ (96)
The region of no propagation shrinks down to a point when $\alpha_{V}$ exceeds a threshold value, given by the equation $bc=9ad$, namely
@@
$\kappa_{\Omega}\simeq\frac{3.0642}{\tau_{H}}\,.$ (98)
All three modes merge at this critical value of $k$ for the given chemical potential.
@@
The wave spectrum is shown as the real and imaginary parts of $\tau_{H}\omega$ on the left and right columns of Fig. 5. The dimensionless chemical potential $\alpha_{V}=\mu_{V}/T$ increases from 00 on the first line to $8$ on the last line. At $\alpha_{V}=0$, ${\rm Re}(\tau_{H}\omega)=0$ for $\tau_{H}\kappa_{\Omega}<\tau_{H}k_{\rm th}\Omega/T=\pi^{2}/12\ln 2\simeq 1.18657$. At $\alpha_{V}=1$, ${\rm Re}(\tau_{H}\omega)$ is nonvanishing for $0<\kappa_{\Omega}<\kappa_{\Omega}^{-}$ and when $\kappa_{\Omega}>\kappa_{\Omega}^{+}$, with an intermediate range of finite extent when there is no propagation. At the critical value $\alpha_{V}=1.4471$, there is a single point when $\tau_{H}\kappa_{\Omega}=3.0642$ and none of the modes propagate. Otherwise, for $\alpha_{V}>1.4471$, all modes are propagating.
As we pointed out in section 3.2, the branch of propagating modes at a smaller wavenumber in Fig. 5 can be thought of as the generalization of the Chiral Vortical Wave previously studied in Jiang:2015cva to fluids possessing an independent helical degree of freedom. On the other hand, the modes propagating at a larger wavenumber in Fig. 5 correspond to the generalization of the Helical Vortical Wave found in Ambrus:2019khr for finite vector chemical potential. As the chemical potential is increased, the modes propagate for any wavelength, and distinguishing between CVW and HVW modes becomes problematic.
@@
$a=0.166(2)\ \text{GeV}\,,\qquad b=0.139(16)\ \text{GeV}\,,\qquad c=0.053(21)\ \text{GeV}\,,$
$d=1.308(28)\ \text{GeV}\,,\qquad f=0.273(8)\ \text{GeV}^{-1}\,.$ (100)
@@ 4 Non-conserved axial and helical charges
$\nabla_{\mu}J^{\mu}_{V}=0\,,\qquad\nabla_{\mu}J^{\mu}_{A}=-\dfrac{\mu_{A}T^{2}}{3\tau_{A}}\,,\qquad\nabla_{\mu}J^{\mu}_{H}=-\dfrac{\mu_{H}T^{2}}{3\tau_{H}}\,.$ (101)
The modes supported by Eqs. (101) can be obtained by solving the characteristic equation, ${\rm det}(T^{-2}\mathbb{M})=0$, which reduces to Eq. (37). In what follows, we solve this equation in various limits, as follows: the small chemical potential in Sec. 4.1; small axial relaxation times $\tau_{A},\tau_{H}\rightarrow 0$ in Sections 4.2 and 4.3; the large wavenumber limit $k\rightarrow\infty$ in Sec. 4.4; and the large chemical potential limit $|\mu_{V}|\gg T$ in Sec. 4.5. We summarize our findings in Sec. 4.6
@@ 4.1 Small chemical potential
The small chemical potential limit of Eq. (37) extends Eq. (50) to
@@
$\omega_{{\mathfrak{a}};2}=\frac{3i}{\pi^{2}\tau_{A}(\tau_{A}-\tau_{H}-\tau_{H}\tau_{A}^{2}k^{2}c_{h}^{2})}\left[(\tau_{A}-\tau_{H})\left(1+\frac{12}{49\pi^{2}}\kappa_{\Omega}^{2}\tau_{A}^{2}\right)\right.\\ \left.+\tau_{H}\frac{48}{\pi^{2}}(\ln 2)^{2}\left(1+\frac{3}{28\pi^{2}}\kappa_{\Omega}^{2}\tau_{A}^{2}\right)\right].$ (105)
Recall that the small chemical potential expansion in the unpolarized QED plasma does not work for small wavenumbers $k$. Now, the situation is slightly different: the expansion is unreliable whenever the denominator appearing in the expression for $\omega_{{\mathfrak{a}};2}$ vanishes, and the small $\alpha_{V}$ expansion breaks down. This happens at the “breakdown” wavenumber,
@@
The breakdown of the small chemical potential expansion was already encountered in Sec. 3.1 for conserved axial charge at the level of the fluctuations (c.f. Eq. (55)), giving $k_{\rm br.}=0$. A closer look at the small wavenumber limit in Sec. 3.2 revealed that the axial modes were propagating in the presence of a finite vector chemical potential $\mu_{V}$ (see Eq. (71)). The propagating nature of the axial modes was not captured in the small $\mu_{V}$ expansion of Sec. 3.2. Analogously, at finite $\tau_{A}$ and $\mu_{V}$, we anticipate that the axial modes may be propagating in the vicinity of $k_{\rm br.}$. This intuition will be confirmed with the explicit computation of the spectrum in Sec. 4.6 (see Fig. 9).
@@ 4.2 Small $\tau_{A}$ limit
We can study the limit where the axial charge dissipates much faster than the helical charge. This is somewhat equivalent to freezing the axial degree of freedom. In this limit, the axial mode dissipates very quickly, according to the dispersion relation:
@@
$\omega^{\pm}_{h;0}\simeq\frac{T^{2}}{12\tau_{H}\sigma^{\omega}_{A}}\left(-i\pm\sqrt{\frac{36B^{2}\sigma^{\omega}_{A}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)}\tau_{H}^{2}\kappa_{\Omega}^{2}-1}\right).$ (108)
Therefore, the vector and helical fluctuations propagate provided that the normalized wavenumber $\kappa_{\Omega}$ exceeds some critical value satisfying
@@
The corresponding threshold wavenumber ranges from $k_{*}=T\kappa_{\Omega;*}/(\tau_{H}\Omega)=1/(2\tau_{H}c_{h})$ at vanishing chemical potential [recovering Eq. (52)] to $\tau_{H}\kappa_{\Omega;*}=\pi/4\sqrt{3}\simeq 0.45$ at infinite chemical potential. Note that propagation is inhibited for small wavenumbers at any value of the vector chemical potential, contrary to the situation found for conserved axial charge ($\tau_{A}\to\infty$) in Sec. 3.4 (see Eq. (93) and Fig. 4) where any non-zero chemical potential gave rise to propagating modes in the infrared region with large wavelengths.
In the case of conserved helicity, when $\tau_{H}\to\infty$, the helical vortical modes are propagating for any wavelength, and their dispersion relation Eq. (108) reduces to
$\lim_{(\tau_{A},\tau_{H})\to(0,\infty)}\omega_{h:0}^{\pm}=\pm\frac{\kappa_{\Omega}BT^{2}}{2}\sqrt{\frac{1}{H\sigma^{\omega}_{A}(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)}}\,.$ (110)
@@
$\lim_{(\tau_{A},\tau_{H})\to(0,\infty)}\omega_{h:0}^{\pm}(\mu_{V}\ll T)=\pm c_{h}k\mp 0.0978\dfrac{k\Omega}{T}\alpha_{V}^{2}+O(\alpha_{V}^{4})\,,$ (111)
$\lim_{(\tau_{A},\tau_{H})\to(\infty,\infty)}\omega_{h:0}^{\pm}(\mu_{V}\ll T)=\pm c_{h}k\mp 0.0124\dfrac{k\Omega}{T}\alpha_{V}^{2}+O(\alpha_{V}^{4})\,,$ (112)
where we have used the expansions given in Eq. (32) and the definitions for $\kappa_{\Omega}$ and $c_{h}$ given in Eqs. (28) and(51). In the large chemical potential limit, we can employ the relations (33) to show that both dispersion relations agree quantitatively up to exponentially small corrections. In Fig. 6, we show the propagation velocity $v^{\pm}_{h}\equiv\omega_{h;0}^{\pm}/k$ for conserved helicity in both cases, i.e. conserved axial charge $\tau_{A}\to\infty$ and frozen axial degree of freedom $\tau_{A}\to 0$. Notably, the freezing of the axial degree of freedom results in a diminishing of the propagation velocity at finite vector chemical potential as opposed to the case where chirality is present and conserved.

### v2

@@ 3.3 Large vector chemical potential limit
$\delta\bar{\mu}_{\chi}(t,z)$ $=\frac{\lvert\alpha_{V}\rvert}{\sqrt{3\pi}}\frac{\delta\mu_{V}}{2}\left(e^{-(z-\tilde{v}t)/2\sigma^{2}}-e^{-(z+\tilde{v}t)/2\sigma^{2}}\right),$ (81)
while $\delta\bar{\mu}_{\tilde{\chi}}(t,z)\simeq\delta\bar{\mu}_{\chi}(t,z)$. As shown in Fig. 3, the above solution provides a good match to the numerical solution in the considered limit at the normalized vector chemical potential $\alpha_{V}=\mu_{V}/T=20$. The amplitude of $\delta\bar{\mu}_{\tilde{\chi}}=\delta\bar{\mu}_{A}-s_{V}\delta\bar{\mu}_{H}$ is finite, contrary to the degenerate limit of the case when the charge is conserved when the amplitude of this mode increases like $\alpha_{V}^{4}$ (see Fig. 5 and Eq. (149) from Ref. Morales-Tejera:2024uzg ). To demonstrate the regularization effect of $\tau_{H}$, panel (d) shows the decrease of the ratio $\delta\bar{\mu}_{\tilde{\chi}}/\delta\mu_{V}$ from $O(10^{3})$ at $\tau_{H}\rightarrow\infty$ to $O(1)$ at $c_{h}\tau_{H}/\sigma=10^{6}$. This effect becomes apparent when $\tau_{H}$ is decreased below a value of the order of $e^{\lvert\alpha_{V}\rvert}$.
@@
$\tau_{H}^{-1}(\mu_{V}\gg T)$ $=\frac{\tau^{-1}_{\textrm{deg.}}}{\lvert\alpha_{V}\rvert}e^{-\lvert\alpha_{V}\rvert}\,,$ (82)
with $\tau_{\textrm{deg.}}\simeq 8.432\times 10^{-3}\,{\rm fm}/c$. Taking this dependence into account, we can take the degenerate limit and the helicity conservation limit in a correlated manner. While the propagating modes $\omega_{\pm}$ remain unaffected, the non-propagating mode $\omega_{\textrm{im.}}$ becomes, to leading order,
$\omega_{\textrm{im.}}=-\dfrac{i\pi^{2}}{24\tau_{\textrm{deg.}}\lvert\alpha_{V}\rvert}\,.$ (83)
Therefore, the dissipation is inhibited as we increase the chemical potential, contrary to the naive $\alpha_{V}\to\infty$ limit in Eq. (76). Finally, taking the limit for the amplitudes (3.3) gives
$\delta\mu_{\chi}^{\textrm{im.}}$ $\simeq-\dfrac{2}{\alpha_{V}^{2}}e^{-\lvert\alpha_{V}\rvert}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}\,,$
$\delta\mu_{V}^{\textrm{im.}}$ $\simeq-\dfrac{72i\kappa_{\Omega}\tau_{\textrm{deg}}}{\pi^{2}\lvert\alpha_{V}\rvert}e^{-\lvert\alpha_{V}\rvert}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}.$ (84)
We conclude that the fluctuation amplitudes $\delta\mu_{\chi}^{\rm im}$ and $\delta\mu_{V}^{\rm im}$ are exponentially suppressed compared to the amplitude $\delta\mu^{\rm im}_{\tilde{\chi}}$. In the above limit, we obtain $\delta\mu_{V}^{\pm}\simeq\delta\mu_{V}/2$ and $\delta\mu_{\tilde{\chi}}^{\rm im}\simeq\delta\mu_{\tilde{\chi}}-\delta\mu_{\chi}$ and thus none of the modes have exponentially-large amplitude. The amplitude of the propagating modes $\delta\mu^{\pm}_{\chi}$ and $\delta\mu^{\pm}_{\tilde{\chi}}$ still grows linearly with $\lvert\alpha_{V}\rvert$ [see Eq. (78)], substantially milder than the quartic dependence $\delta\mu^{h;\pm}_{\tilde{\chi}}\sim\alpha_{V}^{4}$ found when the helicity charge is strictly conserved [see Eq. (149) in Ref. Morales-Tejera:2024uzg ].
@@ 3.4 Generic properties of spectra
In Subsec. 3.1 we have seen that, at small vector chemical potential, there is a critical wavenumber above which two of the three modes start to propagate. At the same time, for small wavenumbers at finite chemical potential, we showed in Subsec. 3.2 that again, there are two modes that always propagate. These two results are compatible with each other only if, for small chemical potential, the modes that propagate at small $k$ turn into non-propagating modes at some finite $k$. We will shortly see that this is indeed the case.
Figure 4: Propagating points $\kappa_{\Omega}^{+}$ (upper curve) and $\kappa_{\Omega}^{-}$ (lower curve) as functions of the normalized vector chemical potential $\alpha_{V}=\mu_{V}/T$ for the unpolarized plasma. In the shaded region, there are no propagating modes. The boundaries $\kappa_{\Omega}=\kappa_{\Omega}^{\pm}$ of the shaded region, defined in Eq. (90), join each other at the critical value $\kappa_{\Omega}^{\rm crit.}$ given in Eq. (95). The vertical lines mark the energy of the collision $\sqrt{s}$ corresponding to the given value of the normalized chemical potential $\alpha_{V}$ [see a discussion around Eq. (96)].
@@
$4ac^{3}(\kappa_{\Omega}\tau_{H})^{6}-(b^{2}c^{2}+18abcd-27a^{2}d^{2})(\kappa_{\Omega}\tau_{H})^{4}$
$+4b^{3}d(\tau_{H}\kappa_{\Omega})^{2}=0.$ (89)
@@
$(\kappa_{\Omega}^{\pm})^{2}$ $=\frac{1}{8ac^{3}\tau_{H}^{2}}\left[b^{2}c^{2}+18abcd-27a^{2}d^{2}\right.$
±$\left.\pm\sqrt{bc-ad}(bc-9ad)^{3/2}\right].$ (90)
@@
$\kappa_{\Omega}^{-}=\dfrac{\pi^{2}}{21\tau_{H}\ln^{2}2}\alpha_{V}+O(\alpha_{V}^{3}),$
$\kappa_{\Omega}^{+}=\frac{\pi^{2}}{12\tau_{H}\ln 2}+\dfrac{1.135}{\tau_{H}}\alpha_{V}^{2}+O(\alpha_{V}^{4}).$ (93)
The region of no propagation shrinks down to a point when $\alpha_{V}$ reaches a threshold value, given by the equation $bc=9ad$, namely
@@
$\kappa_{\Omega}^{\rm crit}\simeq\frac{3.0642}{\tau_{H}}\,.$ (95)
All three modes merge at this critical value of $\kappa_{\Omega}$ for the given chemical potential.
@@
The wave spectrum is shown as the real and imaginary parts of $\tau_{H}\omega$ on the left and right columns of Fig. 5. The dimensionless chemical potential $\alpha_{V}=\mu_{V}/T$ increases from 00 on the first line to $8$ on the last line. At $\alpha_{V}=0$, ${\rm Re}(\tau_{H}\omega)=0$ for $\tau_{H}\kappa_{\Omega}<\tau_{H}k_{\rm th}\Omega/T=\pi^{2}/12\ln 2\simeq 1.18657$. At $\alpha_{V}=1$, ${\rm Re}(\tau_{H}\omega)$ is nonvanishing for $0<\kappa_{\Omega}<\kappa_{\Omega}^{-}\simeq 1.51\tau_{H}^{-1}$ and when $\kappa_{\Omega}>\kappa_{\Omega}^{+}\simeq 1.88\tau_{H}^{-1}$, with an intermediate range of finite extent where there is no propagation. At the critical value $\alpha_{V}=1.4471$, this range is reduced to a single point, $\tau_{H}\kappa_{\Omega}^{+}=\tau_{H}\kappa_{\Omega}^{-}\equiv\tau_{H}\kappa_{\Omega}^{\rm crit}=3.0642$. Otherwise, for $\alpha_{V}>1.4471$, all modes are propagating.
As we pointed out in section 3.2, the branch of propagating modes at a smaller wavenumber in Fig. 5 can be thought of as the generalization of the Chiral Vortical Wave previously studied in Jiang:2015cva ; Gorbar:2017toh to fluids possessing an independent helical degree of freedom. On the other hand, the modes propagating at a larger wavenumber in Fig. 5 correspond to the generalization of the Helical Vortical Wave found in Ambrus:2019khr for finite vector chemical potential. As the chemical potential is increased, the modes propagate for any wavelength, and distinguishing between CVW and HVW modes becomes problematic.
@@
$a=0.166(2)\ \text{GeV}\,,\qquad b=0.139(16)\ \text{GeV}\,,$
$c=0.053(21)\ \text{GeV}\,,\qquad d=1.308(28)\ \text{GeV}\,,$
$f=0.273(8)\ \text{GeV}^{-1}\,.$ (97)
@@ 4 Non-conserved axial and helical charges
$\partial_{\mu}J^{\mu}_{V}=0\,,\ \partial_{\mu}J^{\mu}_{A}=-\dfrac{\mu_{A}T^{2}}{3\tau_{A}}\,,\ \partial_{\mu}J^{\mu}_{H}=-\dfrac{\mu_{H}T^{2}}{3\tau_{H}}\,.$ (98)
The modes supported by Eqs. (98) can be obtained by solving the characteristic equation, ${\rm det}(T^{-2}\mathbb{M})=0$, which reduces to Eq. (33). In what follows, we solve this equation in various limits: the small chemical potential in Sec. 4.1; small axial relaxation times $\tau_{A},\tau_{H}\rightarrow 0$ in Sections 4.2 and 4.3; the large wavenumber limit $k\rightarrow\infty$ in Sec. 4.4; and the large chemical potential limit $\lvert\mu_{V}\rvert\gg T$ in Sec. 4.5. We summarize our findings in Sec. 4.6
@@ 4.1 Small chemical potential
Writing $\omega=\omega_{0}+\omega_{1}\alpha_{V}+\omega_{2}\alpha_{V}^{2}+\dots$ as in Eq. (45), the small chemical potential limit of Eq. (33) extends Eq. (46) to
@@
$\omega_{{\mathfrak{a}};2}=\frac{3i}{\pi^{2}\tau_{A}(\tau_{A}-\tau_{H}-\tau_{H}\tau_{A}^{2}k^{2}c_{h}^{2})}\\ \times\left[(\tau_{A}-\tau_{H})\left(1+\frac{12}{49\pi^{2}}\kappa_{\Omega}^{2}\tau_{A}^{2}\right)\right.\\ \left.+\tau_{H}\frac{48}{\pi^{2}}(\ln 2)^{2}\left(1+\frac{3}{28\pi^{2}}\kappa_{\Omega}^{2}\tau_{A}^{2}\right)\right].$ (102)
Recall that the small chemical potential expansion in the conserved axial charge ($\tau_{A}\rightarrow\infty$) case does not work for small wavenumbers $k$. Now, the situation is slightly different: the expansion is unreliable whenever the denominator appearing in the expression for $\omega_{{\mathfrak{a}};2}$ vanishes, and the small $\alpha_{V}$ expansion breaks down. This happens at the “breakdown” wavenumber,
@@
The breakdown of the small chemical potential expansion was already encountered in Sec. 3.1 for conserved axial charge at the level of the fluctuation amplitudes (c.f. Eq. (51)), giving $k_{\rm br.}=0$. A closer look at the small wavenumber limit in Sec. 3.2 revealed that the axial modes were propagating in the presence of a finite vector chemical potential $\mu_{V}$ (see Eq. (67)). The propagating nature of the axial modes was not captured in the small $\mu_{V}$ expansion of Sec. 3.2. Analogously, at finite $\tau_{A}$ and $\mu_{V}$, we anticipate that the axial modes may be propagating in the vicinity of $k_{\rm br.}$. This intuition will be confirmed with the explicit computation of the spectrum in Sec. 4.6 (see Fig. 9).
@@ 4.2 Small $\tau_{A}$ limit
We can study the limit when the axial charge dissipates much faster than the helical charge. This is equivalent to freezing the axial degree of freedom. In this limit, the axial mode dissipates very quickly, according to the dispersion relation:
@@
$\omega^{\pm}_{h;0}\simeq$
$\frac{T^{2}}{12\tau_{H}\sigma^{\omega}_{A}}\left(-i\pm\sqrt{\frac{36B^{2}\sigma^{\omega}_{A}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)}\tau_{H}^{2}\kappa_{\Omega}^{2}-1}\right).$ (105)
Therefore, the vector and helical fluctuations propagate provided that the normalized wavenumber $\kappa_{\Omega}=k\Omega/T$ [see Eq. (15b)] exceeds some critical value satisfying
@@
The corresponding threshold wavenumber ranges from $k_{*}=T\kappa_{\Omega;*}/(\tau_{H}\Omega)=1/(2\tau_{H}c_{h})$ at vanishing chemical potential [recovering Eq. (48)] to a value $k_{*}=T\kappa_{\Omega;*}/\Omega$ satisfying $\tau_{H}\kappa_{\Omega;*}=\pi/4\sqrt{3}\simeq 0.45$ at infinite chemical potential. Note that propagation is inhibited for small wavenumbers at any value of the vector chemical potential, contrary to the situation found for conserved axial charge ($\tau_{A}\to\infty$) in Sec. 3.4 (see Eq. (90) and Fig. 4) where any non-zero chemical potential gave rise to propagating modes in the infrared region with large wavelengths.
In the case of conserved helicity, when $\tau_{H}\to\infty$, the helical vortical modes are propagating for any wavelength, and their dispersion relation Eq. (4.2) reduces to
$\lim_{(\tau_{A},\tau_{H})\to(0,\infty)}\omega_{h:0}^{\pm}=\pm\frac{\kappa_{\Omega}BT^{2}}{2\sqrt{H\sigma^{\omega}_{A}(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)}}\,.$ (107)
@@
$\lim_{\begin{subarray}{c}\tau_{A}\rightarrow 0\\ \tau_{H}\rightarrow\infty\end{subarray}}\omega_{h:0}^{\pm}(\mu_{V}\ll T)$ =±$=\pm c_{h}k\mp 0.0978\dfrac{k\Omega}{T}\alpha_{V}^{2}$
$\hskip 86.72267pt+O(\alpha_{V}^{4})\,,$ (108a)
$\lim_{\begin{subarray}{c}\tau_{A}\rightarrow\infty\\ \tau_{H}\rightarrow\infty\end{subarray}}\omega_{h:0}^{\pm}(\mu_{V}\ll T)$ =±$=\pm c_{h}k\mp 0.0124\dfrac{k\Omega}{T}\alpha_{V}^{2}$
$\hskip 86.72267pt+O(\alpha_{V}^{4})\,,$ (108b)
where we have used the expansions given in Eq. (29) and the definitions for $\kappa_{\Omega}$ and $c_{h}$ given in Eqs. (15b) and (47). In the large chemical potential limit, we can employ the relations (30) to show that both dispersion relations agree quantitatively up to exponentially small corrections. In Fig. 6, we show the propagation velocity $v^{\pm}_{h}\equiv\omega_{h;0}^{\pm}/k$ for conserved helicity in both cases, i.e. conserved axial charge $\tau_{A}\to\infty$ and frozen axial degree of freedom $\tau_{A}\to 0$. Notably, the freezing of the axial degree of freedom results in a diminishing of the propagation velocity at finite vector chemical potential as opposed to the case when chirality is present and conserved.